# A Nonperturbative Toolkit for Quantum Gravity

### v2

@@ 5.1 A complete basis
In fact, for any topology that supports a saddlepoint, the sum over geometries associated to each topology will match. To see this, consider the contribution of any such geometry. If we fix the geometry consistently with the boundary conditions, in the heavy shell limit the sum over shell trajectories will again localize to the vicinity of spacetime boundary. So the interior gravitational action will be the same, while the shells contribute a universal piece from their trajectory near the fixed boundary.
Similarly we can even include topologies that do not support a saddlepoint, if they exist. To do so we have to show the relevant configurations associated to any of these topologies can we represented as shells propagating on disks with interior topological decorations. We will argue for this below. Then, by the same reasoning as above, provided we take the heavy shell limit first, then for any fixed geometry the shell path integral will universally localize near the boundary, and the $\kappa\to\infty$ limit will restrict the sum to connected geometries. Then the equalities (46) and (48) and (49) follow for the complete gravitational path integral. This argument relies on the heavy shell and $\kappa\to\infty$ limits, which lead to no loss of generality because we are always free to make the basis as overcomplete as we want ($\kappa\to\infty$), and the inertial mass of the shells can be arbitrarily chosen in the basis construction.
@@
into parts that vanish once $n\to-1$, and a remainder that limits to $\overline{\langle\Psi|\Psi\rangle}$. We are using notation of Sec. 2, where $\zeta[\cdots]$ denotes the gravity path integral subject to some boundary conditions and $\phi$ denotes the shell matter fields. We start by organizing the full sum over connected geometries by topology. For any given metric, the path integral for the shell trajectories in the heavy shell limit will localize to geodesics on the background. So for each geometry $\mathcal{M}$ associated to a given topology we can cut out a small region around the $i$-th shell worldvolume of size $\epsilon_{s_{i}}$. There are $n+1$ such regions, one for each shell, see Fig. 20(a). We denote the part of the geometry excluding these regions as $\mathcal{M}_{bulk}$, which contains all handles of $\mathcal{M}$. As the Einstein-Hilbert action consists of local and boundary terms the total action for a given geometry is the sum of the action of these individual patches:
@@
This shows full equality between the path integrals $\overline{\langle\Psi|\prod_{\mathcal{H}_{shell}}|\Psi\rangle}=\overline{\langle\Psi|\Psi\rangle}$. The argument can be extended to show $\overline{(\langle\Psi|\prod_{\mathcal{H}_{shell}}|\Psi\rangle-\langle\Psi|\Psi\rangle)^{2}}=0$ by including all handles connecting $\langle\Psi|\prod_{shell}|\Psi\rangle$ with $\langle\Psi|\prod_{shell}|\Psi\rangle$ (or $\langle\Psi|\Psi\rangle$ in the cross term) into the $I_{bulk}$ region and repeating the above argument. This concludes our argument that the $\kappa\to\infty$ set of shell states at any preparation temperature spans the fine-grained Hilbert space $\mathcal{H}_{LR}$ in the full gravitational path integral.
@@ 5.2.1 Bijection by cutting and gluing
where $I_{local}$ is the local bulk action and $I_{GHY}$ the Gibbons-Hawking-York boundary terms York:1972sj ; Gibbons:1976ue .
@@
The contributions $I_{GHY}$ are determined by the asymptotic boundary structure of the glued wormholes. In gluing sheet diagrams into wormholes the asymptotic boundary sections of the sheet are compactified into circular boundaries as in Fig. 24. To compute the boundary terms the path integral should to be gauge fixed by choosing a conformal frame, say the Fefferman-Graham gauge Fefferman:2007rka for the case of the asymptotically locally AdS gravity path integral. Fixing a frame fixes the asymptotic behavior for any metric in the sum over geometries, and hence the boundary terms are universal and proportional to the length of the asymptotic boundary circle. Hence gluing a sheet into Tr or ZZ wormhole results in the same $I_{GHY}$.
@@
To do this extension we start with a ZZ wormhole geometry $\mathcal{M}_{ZZ}$. We divide the $\mathcal{M}_{ZZ}$ sheet into alternating wedges, each including a certain shell or asymptotic boundary as in Fig. 25(a). For higher topology contributions some of the resulting wedges might be connected by a number of handles. We can take the two wedges containing an identified shell, say the two $c$-wedges on the $L$ sheet and the $d$-wedges on the $R$ sheet, and move one of the $d$ wedges onto the $L$ sheet by swapping it with one of the $c$ wedges, see Fig. 25(b). Doing this for every pair of identified shell we are left with $L$ and $R$ sheets that each have the Tr identification pattern, see Fig. 25(c). The rearrangement produces a Tr geometry from the original ZZ geometry. The smooth metrics on the two spaces will match everywhere except possibly on the sets of measure zero where cut and re-joined them. Hence their actions, computed from local integrals, will be the same. Since these smooth metrics constitute the finite contributions to the path integral we conclude that the path integrals over metrics on the ZZ and Tr geometries should be the same.
@@ 6.1 Algorithm for deriving relational equivalence
It may be that the computation of $\overline{\mathcal{A}}$ requires an analytic continuation $\lim_{n\to-1}\overline{\mathcal{A}}(n)$ involving some insertions of $G^{n}_{ij}$, while this is not the case for $\mathcal{B}$. To show that $\lim_{n\to-1}\overline{\mathcal{A}}(n)=\overline{{\mathcal{B}}}$, reorganize the path integral by using the arguments in Sec. 5.1. Alternatively, as in Sec. 5.2 we can establish equivalence by appropriately inserting a shell resolution of the identity $\mathds{1}_{\mathcal{H}_{LR}}=\lim_{n\to-1}G^{n}_{ij}|i\rangle\langle j|$ in the path integral for $\mathcal{B}$ so that ¯$\overline{\mathcal{B}}=\lim_{n\to-1}\overline{\mathcal{B}}(n)$, where by ¯$\overline{\mathcal{B}}(n)$ we mean the path integral with $\mathds{1}_{\mathcal{H}_{LR}}$ inserted. This procedure allows us to directly compare $\overline{\mathcal{A}}(n)\stackrel{{\scriptstyle?}}{{=}}\overline{\mathcal{B}}(n)$ at fixed positive $n$. We then establish equality by comparing the corresponding sheet diagrams as described in Step 3.
@@
Each of the wormholes contributing to $\overline{\mathcal{A}}(n)$ and ¯$\overline{\mathcal{B}}(n)$ can be cut along the shell worldvolumes to produce a set of sheet diagrams. These serve as a visual tool for organizing the sum over geometries in the $\kappa\to\infty$ limit, and hence are a geometrical analogue of Feynman diagrams. [^2] Note that the sheet diagrams not literally gravitational Feynman diagrams representing graviton scattering. Even if neither $\overline{\mathcal{A}}(n)$ or ¯$\overline{\mathcal{B}}(n)$ can be explicitly calculated, their equivalence can be established by showing a bijection that maps a given $\mathcal{A}$-sheet to a $\mathcal{B}$-sheet with the same action and vice versa.
@@ 6.2 Macroscopic superpositions of semiclassical geometries
In the microcanonical ensemble a black hole saddle can be selected out by imposing a boundary condition fixing the ADM energy (see Chua:2023srl ; Marolf:2018ldl ). Together (38) and (49) imply that microcanonical shell states can also be obtained via the Laplace transform of the canonical states:
[…]
@@
[…]
@@
Hence by choosing $\tilde{\beta}_{L},\tilde{\beta}_{R}>\frac{2\pi\ell}{\sqrt{2}}$, a microcanonical black hole state, associated to a saddlpoint with a horizon, can be written as a superposition of type 2 geometries without horizons, but which have disconnected closed universe components instead (see Fig. 10(b)). Observers inside the compact disconnected universe in the latter cannot send signals to an asymptotic observer, much like observers in the black hole interior. So like the black hole interior the events inside the closed universe component are hidden from infinity. However, unlike for a black hole, the asymptotic observer cannot send a classical probe into this hidden region. That said, the Euclidean state preparation path integral entangles the quantum state in the compact component with the state in the asymptotic component. This means that observers in either region might be able to use the entanglement of the state to send signals to each other, similar to recent proposals that the black hole interior lies within the entanglement wedge of the distant Hawking radiation, at least at late times Almheiri:2019qdq ; Penington:2019kki . See Antonini:2023hdh for a detailed discussion of the encoding of the compact universe into the boundary theory.
@@
See Balasubramanian:2007zt for an example where such a notion can be defined reasonably precisely in the context of the half-BPS states of supersymmetric Yang-Mills and the dual AdS5 gravity. In our case, since the $\tilde{\beta}_{L},\tilde{\beta}_{R}>\frac{2\pi\ell}{\sqrt{2}}$ type 2 shell states form a basis for the entire Hilbert space $\mathcal{H}_{LR}$, and have only a single horizonless saddle computing the norm, we can say that the latter geometric states form a complete basis of eigenvectors for $\mathcal{G}$. But since the microcanonical black hole states also have a single saddle computing the norm, but have a semiclassical geometry with horizons distinct from that of any type 2 state, (66) implies it is impossible for $\mathcal{G}$ to be hermitian with respect to the non-perturbative inner product. A version of this claim was made in Balasubramanian:2022gmo , were it was suggested that micro-canonical type 1 shell with a very long wormhole can be understood as superpositions of short wormholes, suggesting that there is no linear operator that extracts the wormhole length from the underlying state. The nonexistence of a linear geometry operator is compatible with the conjecture that bulk connectedness emerges from quantum entanglement Maldacena:2013xja ; VanRaamsdonk:2010pw , as entanglement cannot be measured by a linear operator.
In the $G_{N}\to 0$ limit, shell states at any temperature are linearly independent, and hence the operator (68) can be Hermitian with respect to the perturbative inner product. This matches expectations from AdS/CFT that the semiclassical gravity description of the CFT emerges in the strict $N\to\infty$ limit, which translates to the $G_{N}\to 0$ limit. Related ideas have been developed from the perspective of state-dependence of the description of black hole interiors Papadodimas:2015jra ; Papadodimas:2015xma , non-isometry in the holographic code Akers:2022qdl ; Antonini:2024yif , the nonexistence of a topology operator Jafferis:2017tiu , the exterior description of a black hole Concepcion:2024nwv and the algebraic description of the semiclassical limit in gravity Leutheusser:2021qhd ; Leutheusser:2022bgi ; Chandrasekaran:2022eqq ; Chandrasekaran:2022cip ; Witten:2021unn .
@@
[…]
@@ 6.3 The gravity Hilbert space in the presence of an observer
To find a way of defining what we mean by saying that a state has a definite geometry, it may help to understand how observers should be included in the gravitational path integral. Recently, Abdalla:2025gzn proposed a rule for calculating observables in the presence of a gravitating observer (see also Harlow:2025pvj ; Akers:2025ahe for discussion of rules for defining observers). The idea is that for every inner product inserted in the gravitational path integral, the observer’s worldline should propagate between the boundaries defining the corresponding bra and ket, and not to any other. Using this rule Abdalla:2025gzn showed that the Hilbert spaces of closed universes and the black hole interior are larger relative to an observer. These results suggest that we should imagine an additional holographic theory living on the world-line of the observer. It would interesting to use the simplified setting provided by the toolkit we have developed to probe this observer Hilbert space in more detail. For example, an observer should be able to probe its local patch of spacetime, and therefore it is unlikely that geometries with a horizon can be expanded in terms of geometries without in the presence of an observer. Such an effect would reduce the linear dependence of geometries shown in our results, perhaps providing a physical mechanism for the increase in the dimension of the Hilbert space when an observer is present.
Recently Antonini:2024mci used AdS/CFT to argue that two distinct semiclassical geometries can be assigned to the type 2 shell states. One is the leading saddle in the gravitational path integral computing its norm, discussed in (3), and the other is obtained by the extrapolate dictionary, which results in two copies of empty AdS with some operator insertion, but without a compact big-crunch universe. It would be interesting to understand whether including an observer in the type 2 shell states and applying the rule from Abdalla:2025gzn can reconcile these two views. Indeed Engelhardt:2025vsp argued that there are asymptotic observables that can detect the presence or absence of a semiclassical compact universe. In our work we have argued that in some circumstances a space with a horizon can be regarded as superposition of the type 2 shell states discussed in Antonini:2024mci and Engelhardt:2025vsp . Our is not in tension with theirs because we are considering a very large quantum superstition of the type 2 states. There is nothing semiclassical about it.

### v3

@@ 5.1 A complete basis
In fact, for any topology that supports a saddlepoint, the sum over geometries associated to each topology will match. To see this, consider the contribution of any such geometry. If we fix the geometry consistently with the boundary conditions, in the heavy shell limit the sum over shell trajectories will again localize to the vicinity of the fixed spacetime boundary. So the interior gravitational action will be the same, while the shells contribute a universal piece from their trajectory near the fixed boundary.
We can also attempt to include topologies that do not support a classical saddle point, if such contributions exist. Below we outline a speculative schematic procedure for how this might be done. As a cautionary note, we do not claim that the path-integral manipulations presented here are strictly well defined, particularly in dimensions $d\geq 4$. Our argument also relies on several heuristic assumptions about the integration measure and the analytic continuation of certain contributions. Although this presentation is non-rigorous, we expect the schematic to capture the essential features of a careful implementation of these ideas in lower-dimensional gravity models, where the relevant analytic and geometric structures are better controlled.
To do so we have to show the relevant configurations associated to any of these topologies can we represented as shells propagating on disks with interior topological decorations. We will argue for this below. Then, by the same reasoning as above, provided we take the heavy shell limit first, then for any fixed geometry the shell path integral will universally localize near the boundary, and the $\kappa\to\infty$ limit will restrict the sum to connected geometries. Then the equalities (47) and (49) and (50) follow for the complete gravitational path integral. This argument relies on the heavy shell and $\kappa\to\infty$ limits, which lead to no loss of generality because we are always free to make the basis as overcomplete as we want ($\kappa\to\infty$), and the inertial mass of the shells can be arbitrarily chosen in the basis construction.
@@
into parts that vanish once $n\to-1$, and a remainder that limits to $\overline{\langle\Psi|\Psi\rangle}$. We are using notation of Sec. 2, where $\zeta[\cdots]$ denotes the gravity path integral subject to some boundary conditions and $\phi$ denotes the shell matter fields.
We start by organizing the full sum over connected geometries by topology. We assume the path integral (56) can be performed by first performing performing the matter path integral on each geometry $g$ contributing at a fixed topology, and subsequently integrating over these geometries. For a given metric, we can then consider the semiclassical limit of the matter path integral such that the shell trajectories in the heavy shell limit localize to geodesics on that background. So for each geometry $\mathcal{M}$ associated to a given topology we can cut out a small region around the $i$-th shell worldvolume of size $\epsilon_{s_{i}}$. There are $n+1$ such regions, one for each shell, see Fig. 20(a). We denote the part of the geometry excluding these regions as $\mathcal{M}_{bulk}$, which contains all handles of $\mathcal{M}$. As the Einstein-Hilbert action consists of local and boundary terms the total action for a given geometry is the sum of the action of these individual patches:
@@
Schematically, this approach establishes full equality between the path integrals, $\overline{\langle\Psi|\prod_{\mathcal{H}_{\text{shell}}}|\Psi\rangle}=\overline{\langle\Psi|\Psi\rangle}$. Although many of the intermediate steps require a more rigorous justification, we expect this argument captures the essential qualitative features of a more complete treatment. Similarly, the argument could be extended to show $\overline{(\langle\Psi|\prod_{\mathcal{H}_{shell}}|\Psi\rangle-\langle\Psi|\Psi\rangle)^{2}}=0$ by including all handles connecting $\langle\Psi|\prod_{shell}|\Psi\rangle$ with $\langle\Psi|\prod_{shell}|\Psi\rangle$ (or $\langle\Psi|\Psi\rangle$ in the cross term) into the $I_{bulk}$ region and repeating the above argument. This concludes our argument that the $\kappa\to\infty$ set of shell states at any preparation temperature spans the fine-grained Hilbert space $\mathcal{H}_{LR}$ in the full gravitational path integral.
@@ 5.2.1 Bijection by cutting and gluing
where $I_{local}$ is the local bulk action and $I_{GHY}$ the Gibbons-Hawking-York boundary terms York (1972); Gibbons and Hawking (1977).
@@
The contributions $I_{GHY}$ are determined by the asymptotic boundary structure of the glued wormholes. In gluing sheet diagrams into wormholes the asymptotic boundary sections of the sheet are compactified into circular boundaries as in Fig. 24. To compute the boundary terms the path integral should to be gauge fixed by choosing a conformal frame, say the Fefferman-Graham gauge Fefferman and Graham (2011) for the case of the asymptotically locally AdS gravity path integral. Fixing a frame fixes the asymptotic behavior for any metric in the sum over geometries, and hence the boundary terms are universal and proportional to the length of the asymptotic boundary circle. Hence gluing a sheet into Tr or ZZ wormhole results in the same $I_{GHY}$.
@@
To do this extension we start with a ZZ wormhole geometry $\mathcal{M}_{ZZ}$. We divide the $\mathcal{M}_{ZZ}$ sheet into alternating wedges, each including a certain shell or asymptotic boundary as in Fig. 25(a). For higher topology contributions some of the resulting wedges might be connected by a number of handles. We can take the two wedges containing an identified shell, say the two $c$-wedges on the $L$ sheet and the $d$-wedges on the $R$ sheet, and move one of the $d$ wedges onto the $L$ sheet by swapping it with one of the $c$ wedges, see Fig. 25(b). Doing this for every pair of identified shell we are left with $L$ and $R$ sheets that each have the Tr identification pattern, see Fig. 25(c). The rearrangement produces a Tr geometry from the original ZZ geometry. The smooth metrics on the two spaces will match everywhere except possibly on the sets of measure zero where cut and re-joined them. Hence their actions, computed from local integrals, will be the same. Since these smooth metrics constitute the finite contributions to the path integral we conclude that the path integrals over metrics on the ZZ and Tr geometries should be the same. Here we assume that the re-glued geometry has the same measure factor as the original geometry. This requires more rigorous justification. It would be interesting to show this explicitly in lower dimensional models of gravity.
@@ 6.1 Algorithm for deriving relational equivalence
It may be that the computation of $\overline{\mathcal{A}}$ requires an analytic continuation $\lim_{n\to-1}\overline{\mathcal{A}}(n)$ involving some insertions of $G^{n}_{ij}$, while this is not the case for $\mathcal{B}$. To show that $\lim_{n\to-1}\overline{\mathcal{A}}(n)=\overline{{\mathcal{B}}}$, reorganize the path integral by using the arguments in Sec. 5.1. Alternatively, as in Sec. 5.2 we can establish equivalence by appropriately inserting a shell resolution of the identity $\mathds{1}_{\mathcal{H}_{LR}}=\lim_{n\to-1}G^{n}_{ij}|i\rangle\langle j|$ in the path integral for $\mathcal{B}$ so that $\overline{\mathcal{B}}=\lim_{n\to-1}\overline{\mathcal{B}}(n)$, where by $\overline{\mathcal{B}}(n)$ we mean the path integral with $\mathds{1}_{\mathcal{H}_{LR}}$ inserted. This procedure allows us to directly compare $\overline{\mathcal{A}}(n)\stackrel{{\scriptstyle?}}{{=}}\overline{\mathcal{B}}(n)$ at fixed positive $n$. We then establish equality by comparing the corresponding sheet diagrams as described in Step 3.
@@
Each of the wormholes contributing to $\overline{\mathcal{A}}(n)$ and $\overline{\mathcal{B}}(n)$ can be cut along the shell worldvolumes to produce a set of sheet diagrams. These serve as a visual tool for organizing the sum over geometries in the $\kappa\to\infty$ limit, and hence are a geometrical analogue of Feynman diagrams. [^6] Note that the sheet diagrams not literally gravitational Feynman diagrams representing graviton scattering. Even if neither $\overline{\mathcal{A}}(n)$ or $\overline{\mathcal{B}}(n)$ can be explicitly calculated, their equivalence can be established by showing a bijection that maps a given $\mathcal{A}$-sheet to a $\mathcal{B}$-sheet with the same action and vice versa. We argued that such equalities can be naturally extended to include perturbative corrections around the saddle points. We also outlined how these equalities might be further generalized to incorporate geometries in the path integral whose topologies do not admit a classical saddle point. Although this outline relies on several heuristic assumptions regarding the integration measure and the analytic continuation of certain contributions, we expect the schematic to capture the essential qualitative features. It would be interesting to realize these ideas explicitly in lower-dimensional models of gravity, such as JT gravity or AdS3 gravity.
@@ 6.2 Macroscopic superpositions of semiclassical geometries
In the microcanonical ensemble a black hole saddle can be selected out by imposing a boundary condition fixing the ADM energy (see Chua and Hartman (2024); Marolf (2018)). Together (39) and (50) imply that microcanonical shell states can also be obtained via the Laplace transform of the canonical states:
$|i\rangle_{E_{L},E_{R}}\propto\int d\beta_{L}d\beta_{R}e^{\frac{\beta_{R}E_{R}+\beta_{L}E_{L}}{2}}\sqrt{Z(\beta_{L})Z(\beta_{R})}|i,\beta_{L},\beta_{R}\rangle.$ (71)
@@
$|i\rangle_{E_{L},E_{R}}\propto\sum_{j}\int d\beta_{L}d\beta_{R}e^{\frac{\beta_{R}E_{R}+\beta_{L}E_{L}}{2}}a_{j}(i,\beta_{L},\beta_{R})\sqrt{Z(\beta_{L})Z(\beta_{R})}|j,\tilde{\beta}_{L},\tilde{\beta}_{R}\rangle\equiv$ (72)
@@
Hence by choosing $\tilde{\beta}_{L},\tilde{\beta}_{R}>\frac{2\pi\ell}{\sqrt{2}}$, a microcanonical black hole state, associated to a saddlpoint with a horizon, can be written as a superposition of type 2 geometries without horizons, but which have disconnected closed universe components instead (see Fig. 10(b)). Observers inside the compact disconnected universe in the latter cannot send signals to an asymptotic observer, much like observers in the black hole interior. So like the black hole interior the events inside the closed universe component are hidden from infinity. However, unlike for a black hole, the asymptotic observer cannot send a classical probe into this hidden region. That said, the Euclidean state preparation path integral entangles the quantum state in the compact component with the state in the asymptotic component. This means that observers in either region might be able to use the entanglement of the state to send signals to each other, similar to recent proposals that the black hole interior lies within the entanglement wedge of the distant Hawking radiation, at least at late times Almheiri et al. (2020); Penington et al. (2022). See Antonini et al. (2023) for a detailed discussion of the encoding of the compact universe into the boundary theory.
@@
See Balasubramanian et al. (2007) for an example where such a notion can be defined reasonably precisely in the context of the half-BPS states of supersymmetric Yang-Mills and the dual AdS5 gravity. In our case, since the $\tilde{\beta}_{L},\tilde{\beta}_{R}>\frac{2\pi\ell}{\sqrt{2}}$ type 2 shell states form a basis for the entire Hilbert space $\mathcal{H}_{LR}$, and have only a single horizonless saddle computing the norm, we can say that the latter geometric states form a complete basis of eigenvectors for $\mathcal{G}$. But since the microcanonical black hole states also have a single saddle computing the norm, but have a semiclassical geometry with horizons distinct from that of any type 2 state, (72) implies it is impossible for $\mathcal{G}$ to be hermitian with respect to the non-perturbative inner product. A version of this claim was made in Balasubramanian et al. (2024c), were it was suggested that micro-canonical type 1 shell with a very long wormhole can be understood as superpositions of short wormholes, suggesting that there is no linear operator that extracts the wormhole length from the underlying state. The nonexistence of a linear geometry operator is compatible with the conjecture that bulk connectedness emerges from quantum entanglement Maldacena and Susskind (2013); Van Raamsdonk (2010), as entanglement cannot be measured by a linear operator.
In the $G_{N}\to 0$ limit, shell states at any temperature are linearly independent, and hence the operator (74) can be Hermitian with respect to the perturbative inner product. This matches expectations from AdS/CFT that the semiclassical gravity description of the CFT emerges in the strict $N\to\infty$ limit, which translates to the $G_{N}\to 0$ limit. Related ideas have been developed from the perspective of state-dependence of the description of black hole interiors Papadodimas and Raju (2016, 2015), non-isometry in the holographic code Akers et al. (2024); Antonini et al. (2025), the nonexistence of a topology operator Jafferis (2017), the exterior description of a black hole Concepcion et al. (2024) and the algebraic description of the semiclassical limit in gravity Leutheusser and Liu (2023, 2025); Chandrasekaran et al. (2023b, a); Witten (2022).
@@
Despite the considerations described above, bulk observers can measure the spacetime they find themselves in. Thus, some meaningful non-perturbative definition of bulk geometry is necessary. In lower dimensional models of gravity a basis of states explicitly associated with bulk time-slice geometries can be constructed. In two dimensional Jackiw-Teitelboim (JT) gravity the perturbative Hilbert space is spanned by length basis $|\ell\rangle$, which are eigenstates of the operator $\hat{\ell}$ that measures the geodesic length between the two asymptotic boundaries (see for example Harlow and Jafferis (2020); Yang (2019)). However, as shown in Iliesiu et al. (2024) these states too are over-complete due to non-perturbative effects and therefore $\hat{\ell}$ is also not a well defined linear operator in the nonpertubative theory. To resolve this they proposed a nonperturbative definition of the length operator: $\mathcal{F}^{-1}\left(\int d\ell\mathcal{F}(\ell)|\ell\rangle\langle\ell|\right)$ for some real and invertible function $\mathcal{F}$. As this operator maps null states to null states it is a well defined linear operator in the quantum theory. Several authors approached this problem from the alternative perspective of a dual Double Scaled SYK (DSSYK) model. First, Berkooz et al. (2018) showed that that the large system, low energy limit, of the DSSYK theory is a classical limit, and has an effective description that matches JT gravity in two dimensions. Then, Lin (2022); Rabinovici et al. (2023) pointed out that the length of wormholes in JT gravity is precisely computed by the spread complexity of Balasubramanian et al. (2022) of the DSSYK model. Finally, Balasubramanian et al. (2024d) showed how, away from the classical limit, small quantum effects truncate the apparently infinite dimensional Hilbert space identified by Berkooz et al. (2018) to finite dimension. For similar reasons even if a bulk time-slice Hilbert space can be constructed in higher dimensions we expect it also to be overcomplete. Hence it again cannot support a linear geometry operator.
@@ 6.3 The gravity Hilbert space in the presence of an observer
To find a way of defining what we mean by saying that a state has a definite geometry, it may help to understand how observers should be included in the gravitational path integral. Recently, Abdalla et al. (2025) proposed a rule for calculating observables in the presence of a gravitating observer (see also Harlow et al. (2025); Akers et al. (2025) for discussion of rules for defining observers). The idea is that for every inner product inserted in the gravitational path integral, the observer’s worldline should propagate between the boundaries defining the corresponding bra and ket, and not to any other. Using this rule Abdalla et al. (2025) showed that the Hilbert spaces of closed universes and the black hole interior are larger relative to an observer. These results suggest that we should imagine an additional holographic theory living on the world-line of the observer. It would interesting to use the simplified setting provided by the toolkit we have developed to probe this observer Hilbert space in more detail. For example, an observer should be able to probe its local patch of spacetime, and therefore it is unlikely that geometries with a horizon can be expanded in terms of geometries without in the presence of an observer. Such an effect would reduce the linear dependence of geometries shown in our results, perhaps providing a physical mechanism for the increase in the dimension of the Hilbert space when an observer is present.
Recently Antonini and Rath (2024) used AdS/CFT to argue that two distinct semiclassical geometries can be assigned to the type 2 shell states. One is the leading saddle in the gravitational path integral computing its norm, discussed in (3), and the other is obtained by the extrapolate dictionary, which results in two copies of empty AdS with some operator insertion, but without a compact big-crunch universe. It would be interesting to understand whether including an observer in the type 2 shell states and applying the rule from Abdalla et al. (2025) can reconcile these two views. Indeed Engelhardt and Gesteau (2025) argued that there are asymptotic observables that can detect the presence or absence of a semiclassical compact universe. In our work we have argued that in some circumstances a space with a horizon can be regarded as superposition of the type 2 shell states discussed in Antonini and Rath (2024) and Engelhardt and Gesteau (2025). Our is not in tension with theirs because we are considering a very large quantum superstition of the type 2 states. There is nothing semiclassical about it.